# Improved Baselines with Synchronized Encoding for Universal Medical Image Segmentation

### v2

@@ Abstract
Large foundation models, known for their strong zero-shot generalization capabilities, can be applied to a wide range of downstream tasks. However, developing foundation models for medical image segmentation poses a significant challenge due to the domain gap between natural and medical images. While fine-tuning techniques based on the Segment Anything Model (SAM) have been explored, they primarily focus on scaling up data or refining inference strategies without incorporating domain-specific architectural designs, limiting their zero-shot performance. To optimize segmentation performance under standard inference settings and provide a strong baseline for future research, we introduce SyncSAM, which employs a synchronized dual-branch encoder that integrates convolution and Transformer features in a synchronized manner to enhance medical image encoding, and a multi-scale dual-branch decoder to preserve image details. SyncSAM is trained on two of the largest medical image segmentation datasets, SA-Med2D-20M and IMed-361M, resulting in a series of pre-trained models for universal medical image segmentation. Experimental results demonstrate that SyncSAM not only achieves state-of-the-art performance on test sets but also exhibits strong zero-shot capabilities on unseen datasets. The code and model weights are available at https://github.com/Hhankyangg/SyncSAM.
@@ 1 Introduction
Large foundation models pre-trained on extensive datasets provide strong zero-shot capabilities, making them effective for diverse downstream tasks [2, 16, 13, 1]. A key example is the Segment Anything Model (SAM) [11], an interactive segmentation model trained on large datasets. Leveraging its robust zero-shot performance, SAM has not only established a stable foundation for segmentation tasks [14, 12] but also advanced various visual and multimodal applications [19].
Medical image segmentation is a fundamental task in medical imaging analysis, playing a crucial role in disease diagnosis, treatment planning, and disease progression monitoring [6]. With the emergence of SAM, a corpus of researchers have explored its potential in medical image segmentation. These works can be categorized based on the dataset scale: leveraging SAM for specialist models on small datasets and for foundation models on large-scale datasets. For specialist models trained on small datasets, approaches include Parameter-Efficient Fine-Tuning (PEFT) [9, 28, 27] and architectural modifications such as hierarchical decoders [5] and prompt-decoupled mask decoders [7]. While effective on small datasets, their performance on large datasets remains unverified. For foundation models trained on large-scale datasets, two primary fine-tuning strategies exist. The first fine-tunes SAM at different parameter scales using either PEFT techniques [3, 4] or full-parameter fine-tuning [17], yet these models often lack domain-specific architectural enhancements, leading to suboptimal performance. The second modifies the inference process [24, 25] to improve segmentation accuracy, incorporating advanced interactive methods or support-set-based strategies. While effective in complex prompt settings, these modifications may limit broader applicability to downstream tasks.
To fill the gap of foundation models in integrating expert-designed modules, we introduce SyncSAM in this work. Inspired by the advantages of Convolutional Neural Network (CNN) in capturing fine-grained features and dealing with noisy images, we incorporate a CNN branch into SAM’s image encoder, injecting medical-specific domain bias into the Vision Transformer(ViT) backbone. Moreover, we propose a synchronized fusion strategy that performs stage-wise fusion of ViT and CNN features rather than merging them in a single step. This progressive alignment of local and global representations enhances contextual understanding of medical images. To further improve segmentation, we devise a multi-scale dual-branch decoder that leverages early-stage encoder features to refine fine-grained edge details. By combining the synchronized dual-branch encoder and the multi-scale dual-branch decoder, SyncSAM effectively enhances feature representation, enabling more precise segmentation for medical images.
@@ 2.1 Overview
As illustrated in Fig. 1(a), SyncSAM enhances medical image segmentation with a synchronized dual-branch architecture. Given a pre-processed medical image $I\in\mathbb{R}^{3\times H\times W}$ and a prompt $c$ (points, boxes, or masks), SyncSAM predicts the corresponding binary mask $M^{(c)}$:
@@
SyncSAM consists of three main components: a synchronized dual-branch encoder, a prompt encoder, and a multi-scale dual-branch decoder. The synchronized dual-branch encoder retains SAM’s ViT while integrating a CNN-based synchronized fusion branch to align features at each stage. The prompt encoder, inherited from SAM, converts sparse and dense prompts into vector representations. The decoder then integrates feature maps with prompt information, while the multi-scale fusion branch enhances feature refinement. Together with the proposed Med-Output Token, these components collaboratively produce the final segmentation output. The following sections detail each component.
@@ 2.2 Synchronized Dual-Branch Encoder
Full fine-tuning of SAM on large-scale medical datasets is both computationally expensive and suboptimal (Sec. 3.3). Instead, we adopt partial fine-tuning, since existing PEFT methods lack domain-specific designs. To balance generalization and domain-specific feature extraction, we introduce a dual-branch encoder that combines SAM’s frozen ViT branch with a CNN branch for fine-grained feature capture in noisy images. We further propose a synchronized fusion strategy that fuses features at each encoding stage (Fig. 1(c)); ablation studies (Sec. 3.5) confirm its superiority over single-step fusion.
The synchronized dual-branch encoder, serving as the image encoder in a SAM-like model, encodes a preprocessed image $I\in\mathbb{R}^{3\times H\times W}$ into a rich semantic feature map $F\in\mathbb{R}^{256\times\frac{H}{16}\times\frac{W}{16}}$. Initially, the input image undergoes patch embedding in the ViT branch and stage 0 processing in the CNN branch as a preprocessing step. The image is then processed through four stages in both branches. As shown in Fig. 1(c), at each stage $i$, the ViT branch applies windowed attention transformer blocks, updating the previous stage’s output feature map $F_{\text{output}}^{(i-1)}$ to $F_{\text{windowed}}^{(i)}$, while the CNN branch extracts local features from $F_{\text{CNN}}^{(i-1)}$, producing $F_{\text{CNN}}^{(i)}$. The CNN features $F_{\text{CNN}}^{(i)}$ are then convolved to match the ViT feature shape, yielding $\tilde{F}_{\text{CNN}}^{(i)}$, which is point-wise added to $F_{\text{windowed}}^{(i)}$ to form the fused feature map $F_{\text{fused}}^{(i)}$. The global attention transformer block processes $F_{\text{fused}}^{(i)}$, generating the stage output $F_{\text{output}}^{(i)}$. Finally, the ViT neck layer refines the output, producing the final encoded feature map $F$.
@@ 2.3 Multi-Scale Dual-Branch Decoder
Multi-Scale Fusion Branch. A multi-scale fusion branch is incorporated into the mask decoder, as early-stage features preserve more fine-grained edge details. This branch fuses feature maps $F_{i}$ from the first three ViT stages with a dimensionally adjusted $F$ from the image encoder, generating a multi-scale feature map $\tilde{F}\in\mathbb{R}^{32\times\frac{H}{4}\times\frac{W}{4}}$. The feature fusion is performed via a series of convolutional layers. $\tilde{F}$ is then point-wise added to the mask feature $F_{M}$, which is derived from $F$ using a Two-Way Transformer to integrate information from the output tokens, including prompt tokens and the Med-Output Token.
@@ 3.1 Training of SyncSAM
Following previous studies [11, 17, 4], we train SyncSAM using interactive segmentation simulation, and randomly select five masks per image for a training step. We empirically set $\lambda=20$. For preprocessing, images are padded to square shapes and resized to $256\times 256$ before being fed into the model. We use the ViT-B version of SAM, keeping its image encoder frozen while integrating different CNN branches—ResNet50 and ResNet34 [8]. The corresponding models, SyncSAM-50 and SyncSAM-34, contain approximately 138M and 116M parameters, respectively, with only a subset being trainable. Training is conducted for 12 epochs with an initial learning rate of 0.0001. We use 8 NVIDIA Tesla A100 GPUs, processing 50 images and 250 masks per GPU per step.
@@ 3.2 Experimental Setup
We assess SyncSAM through three key experiments: (1) test set evaluations against models trained on the same dataset (Sec. 3.3), (2) zero-shot segmentation performance on external datasets (Sec. 3.4), and (3) ablation studies analyzing the impact of individual components (Sec. 3.5). In the following tables, “SAM” denotes direct inference with SAM, while “FT-SAM” refers to SAM with only the mask decoder trained. All results are reported as Dice Similarity Coefficient (DSC) percentages, with best and second best performances highlighted.
@@ 3.5 Ablation study
We conduct ablation experiments on SA-Med2D-20M [26], using bounding boxes as prompts to evaluate the impact of each module in SyncSAM, as shown in Tab. 5. Rows 1 and 3 demonstrate that incorporating the synchronized dual-branch encoder alone improves the DSC score by 18.9%. Removing the ViT branch (rows 4, 6, and 7) results in inferior performance compared to the dual-branch configuration, confirming the necessity of ViT. Rows 10 and 11 compare different fusion strategies, showing that synchronized fusion improves DSC by 3.3%, validating its effectiveness. Additionally, we assess the impact of decoder modifications across different configurations (rows 3 vs. 8 vs. 11, rows 4 vs. 6 vs. 7, and rows 2 vs. 5), all of which indicate performance gains from decoder adjustments. Finally, regarding the choice of CNN branch, row 9 vs. row 11 shows that replacing ResNet-34 with ResNet-50 increases DSC by 1.2%, demonstrating the model’s scalability with parameter size.
Table 5: Ablation study. $\triangle$ represents single-step fusion; $\bigstar$ for synchronized fusion.
Model Decoder trained with Image Encoder DSC
Multi-scale Med-token ViT (fixed) ResNet-34 ResNet-50 Fusion Test set
SAM $\checkmark$ 66.6
FT-SAM $\checkmark$ 74.6
$\checkmark$ $\checkmark$ $\bigstar$ 85.5
$\checkmark$ 79.1
$\checkmark$ $\checkmark$ $\checkmark$ 78.2
$\checkmark$ $\checkmark$ 82.7
SyncSAM $\checkmark$ $\checkmark$ $\checkmark$ 84.6
$\checkmark$ $\checkmark$ $\checkmark$ $\bigstar$ 86.6
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\bigstar$ 87.0
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\triangle$ 84.9
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\bigstar$ 88.2

### v4

@@ Abstract
Large foundation models, known for their strong zero-shot generalization capabilities, can be applied to a wide range of downstream tasks. However, developing foundation models for medical image segmentation poses a significant challenge due to the domain gap between natural and medical images. While fine-tuning techniques based on the Segment Anything Model (SAM) have been explored, they primarily focus on scaling up data or refining inference strategies without incorporating domain-specific architectural designs, limiting their zero-shot performance. To optimize segmentation performance under standard inference settings and provide a strong baseline for future research, we introduce SyncSAM, which employs a synchronized dual-branch encoder that integrates convolution and Transformer features in a synchronized manner to enhance medical image encoding, and a multi-scale dual-branch decoder to preserve image details. SyncSAM is trained on two of the largest medical image segmentation datasets, SA-Med2D-20M and IMed-361M, resulting in a series of pre-trained models for universal medical image segmentation. Experimental results demonstrate that SyncSAM not only achieves state-of-the-art performance on test sets but also exhibits strong zero-shot capabilities on unseen datasets. Code and checkpoints are available at https://github.com/Hhankyangg/SyncSAM.
@@ 1 Introduction
Large foundation models pre-trained on extensive datasets provide strong zero-shot capabilities, making them effective for diverse downstream tasks [2, 17, 3, 4]. A key example is the Segment Anything Model (SAM) [13], an interactive segmentation model trained on large datasets. Leveraging its robust zero-shot performance, SAM has not only established a stable foundation for segmentation tasks [20] but also advanced various visual and multimodal applications [15, 1].
Medical image segmentation is a fundamental task in medical imaging analysis, playing a crucial role in disease diagnosis, treatment planning, and disease progression monitoring [8]. With the emergence of SAM, a corpus of researchers have explored its potential in medical image segmentation. These works can be categorized based on the dataset scale: leveraging SAM for specialist models on small datasets and for foundation models on large-scale datasets. For specialist models trained on small datasets, approaches include Parameter-Efficient Fine-Tuning (PEFT) [27] and architectural modifications such as hierarchical decoders [7] and prompt-decoupled mask decoders [10]. While effective on small datasets, their performance on large datasets remains unverified. For foundation models trained on large-scale datasets, two primary fine-tuning strategies exist. The first fine-tunes SAM at different parameter scales using either PEFT techniques [6] or full-parameter fine-tuning [18], yet these models often lack domain-specific architectural enhancements, leading to suboptimal performance. The second modifies the inference process [30, 28] to improve segmentation accuracy, incorporating advanced interactive methods or support-set-based strategies. While effective in complex prompt settings, these modifications may limit broader applicability to downstream tasks [14].
To fill the gap of foundation models in integrating expert-designed modules, we introduce SyncSAM in this work. Inspired by the advantages of Convolutional Neural Network (CNN) in capturing fine-grained features and dealing with noisy images, we incorporate a CNN branch into SAM’s image encoder, injecting medical-specific domain bias into the Vision Transformer(ViT) [9] backbone. Moreover, we propose a synchronized fusion strategy that performs stage-wise fusion of ViT and CNN features rather than merging them in a single step. This progressive alignment of local and global representations enhances contextual understanding of medical images. To further improve segmentation, we devise a multi-scale dual-branch decoder that leverages early-stage encoder features to refine fine-grained edge details. By combining the synchronized dual-branch encoder and the multi-scale dual-branch decoder, SyncSAM effectively enhances feature representation, enabling more precise segmentation for medical images.
@@ 2.1 Overview
Figure. 1(a) illustrates the framework of our proposed SyncSAM, which aims to enhance medical image segmentation with a synchronized dual-branch architecture. Given a pre-processed medical image $I\in\mathbb{R}^{3\times H\times W}$ and a prompt $c$ (points, boxes, or masks), SyncSAM predicts the corresponding binary mask $M^{(c)}$:
@@
SyncSAM consists of three main components: a synchronized dual-branch encoder, a prompt encoder, and a multi-scale dual-branch decoder. The synchronized dual-branch encoder retains SAM’s ViT while integrating a CNN-based synchronized fusion branch to align features at each stage [shown in Fig. 1(b)]. The prompt encoder, inherited from SAM, transforms sparse and dense prompts into vector representations, referred to as prompt token. The decoder then integrates feature maps with prompt token, while the multi-scale fusion branch [detailed in Fig. 1(c)] enhances feature refinement. Together with the proposed Med-Output Token, these components collaboratively produce the final segmentation output. The following sections detail each component.
@@ 2.2 Synchronized Dual-Branch Encoder
The synchronized dual-branch encoder serves as the image encoder in the model, encoding a preprocessed image $I\in\mathbb{R}^{3\times H\times W}$ into a rich semantic feature map $F\in\mathbb{R}^{256\times\frac{H}{16}\times\frac{W}{16}}$. To be specific, the encoder consists of two parallel branches, i.e., a ViT branch and a CNN branch. The ViT branch, inherited from SAM, captures long-range dependencies and global context while remaining frozen to limit the number of trainable parameters. In contrast, the CNN branch extracts local fine-grained features, which are particularly advantageous in handling noise and structural variations commonly found in medical images. Initially, the input image undergoes patch embedding in the ViT branch and stage 0 processing in the CNN branch as a preprocessing step. The image is then progressively processed through four consecutive stages in both branches.
To effectively integrate features from both branches, we design a synchronized fusion mechanism that merges encoded representations at each stage. At stage $i$, the ViT branch leverages windowed attention transformer blocks to refine the output feature map $F_{\text{output}}^{(i-1)}$ from the previous stage, producing $F_{\text{windowed}}^{(i)}$. Simultaneously, the CNN branch extracts local and domain-specific features from $F_{\text{CNN}}^{(i-1)}$, generating $F_{\text{CNN}}^{(i)}$. To ensure compatibility between the two feature maps, $F_{\text{CNN}}^{(i)}$ undergoes convolution to match the ViT feature dimensions, yielding $\tilde{F}_{\text{CNN}}^{(i)}$. This transformed CNN feature map is then point-wise added to $F_{\text{windowed}}^{(i)}$, forming the fused feature map $F_{\text{fused}}^{(i)}$. The fused representation is further refined through a global attention transformer block, producing the stage output $F_{\text{output}}^{(i)}$. Going through all four stages, the encoded feature map is further refined by the ViT neck layer, generating the final output $F$.
@@ 2.3 Multi-Scale Dual-Branch Decoder
Multi-Scale Fusion Branch. A multi-scale fusion branch is incorporated into the mask decoder, as early-stage features preserve more fine-grained edge details [22]. This branch fuses feature maps $F_{i}$ from the first three ViT stages with a dimensionally adjusted $F$ from the image encoder, generating a multi-scale feature map $\tilde{F}\in\mathbb{R}^{32\times\frac{H}{4}\times\frac{W}{4}}$. The feature fusion is performed via a series of convolutional layers. $\tilde{F}$ is then point-wise added to the mask feature $F_{M}$, which is derived from $F$ using a Two-Way Transformer to integrate information from the output tokens, including prompt tokens and the Med-Output Token.
@@ 3.1 Training of SyncSAM
Following previous studies [13, 18, 6], we train SyncSAM using interactive segmentation simulation, and randomly select five masks per image for a training step. We empirically set $\lambda=20$. For preprocessing, images are padded to square shapes and resized to $256\times 256$ before being fed into the model. We use the ViT-B version of SAM, keeping its image encoder frozen while integrating different CNN branches—ResNet50 and ResNet34 [11]. The corresponding models, SyncSAM-50 and SyncSAM-34, contain approximately 138 million and 116 million parameters, respectively, of which about 48 million and 26 million are trainable. Training is conducted for 12 epochs with an initial learning rate of 0.0001. We use 8 NVIDIA Tesla A100 GPUs, processing 50 images and 250 masks per GPU per step.
@@ 3.2 Experimental Setup
We assess SyncSAM through three key experiments: (1) test set evaluations against models trained on the same dataset (Sec. 3.3), (2) zero-shot segmentation performance on external datasets (Sec. 3.4), and (3) ablation studies analyzing the impact of individual components (Sec. 3.5). In the following tables, “SAM” denotes direct inference with SAM, while “FT-SAM” refers to SAM with only the mask decoder trained. All results are reported as Dice Similarity Coefficient (DSC) percentages, with best and second best performances highlighted. All reported scores are averaged over three independent runs.
@@ 3.5 Ablation study
We conduct ablation experiments on SA-Med2D-20M [29], using bounding boxes as prompts to evaluate the impact of each module in SyncSAM, as shown in Tab. 5. Overall, fully fine-tuned SAM achieves a DSC of 80.8 (row 3), while our synchronized encoding variant reaches 85.5 (row 4), highlighting the advantage of our architectural design. Specifically, rows 1 and 4 demonstrate that incorporating the synchronized dual-branch encoder alone improves DSC by 18.9%. Removing the ViT branch (rows 5, 7, and 8) leads to inferior performance compared to the dual-branch configuration, confirming the necessity of retaining ViT for contextual representation. Rows 11 and 12 compare different fusion strategies, showing that synchronized fusion improves DSC by 3.3%, validating its effectiveness. Furthermore, we assess the impact of decoder modifications across various settings (rows 4 vs. 9 vs. 12, rows 5 vs. 7 vs. 8, and rows 2 vs. 6), all of which demonstrate consistent performance gains from decoder enhancements. Lastly, comparing row 10 and row 12 shows that replacing ResNet-34 with ResNet-50 improves DSC by 1.2%, suggesting that our model benefits from increased capacity and is scalable with larger CNN backbones.
Table 5: Ablation. $\triangle$ = single-step fusion; $\bigstar$ = synchronized fusion; RN = ResNet.
# Model Decoder trained with Image Encoder DSC
Multi-scale Med-token ViT (fixed) RN-34 RN-50 Fusion Test set
1 SAM (fixed) $\checkmark$ 66.6
2 FT-SAM $\checkmark$ 74.6
3 SAM 80.8
4 $\checkmark$ $\checkmark$ $\bigstar$ 85.5
5 $\checkmark$ 79.1
6 $\checkmark$ $\checkmark$ $\checkmark$ 78.2
7 $\checkmark$ $\checkmark$ 82.7
8 SyncSAM $\checkmark$ $\checkmark$ $\checkmark$ 84.6
9 $\checkmark$ $\checkmark$ $\checkmark$ $\bigstar$ 86.6
10 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\bigstar$ 87.0
11 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\triangle$ 84.9
12 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\bigstar$ 88.2
@@ 4 Conclusion
{credits}
4.0.1 Acknowledgements
This work was supported by NSFC under Grant 12326615, 42201394 and 12426313, and Key R&D Program of Shaanxi Province under Grant 2025CY-YBXM-040.
[…]
The authors have no competing interests to declare that are relevant to the content of this article.